# DiffPortrait3D: Controllable Diffusion for Zero-Shot Portrait View Synthesis

### 22 Dec 2023 v3

@@ 2 Related Works
GANs [16] employ adversarial learning to synthesize images that faithfully capture the distribution of the training dataset. Previous studies have demonstrated the effectiveness of 2D GANs in portrait manipulation, employing techniques such as latent space exploration [9] and exemplar image utilization [26]. Nevertheless, the absence of inherent 3D representations in these 2D GANs presents a challenge in maintaining 3D consistency for the task of NVS.
@@ Conclusion.
Future work.
While the image coherence is largely strengthened with our cross-view module and 3D-aware generation, we still observe occasionally flickering artifacts in unobserved regions. We leave the exploration of longer-range consistent view manipulation as future work. In this work, the appearance is formulated to be sourced from the reference images only. This could result in some loss of identity given the limited appearance context. In the future, we would like to extend our framework such that the identity can be multi-sourced from e.g., text and personalized Loras [24].
@@
In this supplementary paper, we provide additional implementation details in Section A, showcase more visual results and numerical comparisons in Section B, and discuss limitations in Section C.
@@ B.1 Alignments
B.2 View-consistent novel view synthesis
We show more challenging results in Figure 1314 and  15. Our model is able to generalize well to arbitrary face portraits with unposed camera views, extreme facial expressions, and diverse artistic depictions. Please also refer to our supplementary video for more high-resolution results.
@@ Appendix C Limitation and Future Work
In addition to the limitations and future directions discussed in our main paper, we also include visualizations of failure cases in Figure 12. Rows (a) through (f) display artifacts in areas not observed, accompanied by changes in identity, particularly noticeable in (b), (c), and (d). In cases (a), (e), and (f), it is evident that the model struggles to accurately replicate secondary elements from the reference image, such as sunglasses in (f), hands and flowers in (e), and leaves in (a), leading to inconsistent outcomes. One potential issue we’ve identified stems from the use of a 2D diffusion backbone that integrates 3D-Aware information. This approach, while innovative, may lead to minor inconsistencies and deviations, especially in challenging depictions. Addressing these limitations is an important area that should be addressed in future work.

### 19 Mar 2024 v4

@@ 2 Related Works
GANs [15] employ adversarial learning to synthesize images that faithfully capture the distribution of the training dataset. Previous studies have demonstrated the effectiveness of 2D GANs in portrait manipulation, employing techniques such as latent space exploration [8] and exemplar image utilization [25, 53]. Nevertheless, the absence of inherent 3D representations in these 2D GANs presents a challenge in maintaining 3D consistency for the task of NVS.
@@ Conclusion.
[…]
@@
In this supplementary paper, we provide additional implementation details in Section A, showcase more visual results and numerical comparisons in Section B, and discuss limitations & ethics consideration in Section C and Section  D .
@@ Appendix A Implementation Detail
Figure 11: Ablation on view conditional images.
Figure 12: Novel view synthesis of anthropomorphic animals.
@@ B.1 Alignments
B.2 PanoHead-PTI Comparison
[…]
Figure 14: Novel view synthesis with PanoHead-PTI.
@@ B.2 PanoHead-PTI Comparison
B.3 View-consistent novel view synthesis
We show more challenging results in Figure 1617 and  18. Our model is able to generalize well to arbitrary face portraits with unposed camera views, extreme facial expressions, and diverse artistic depictions. Please also refer to our supplementary video for more high-resolution results.
B.4 Ablation on view condition images
[…]
B.5 Anthropomorphic animals
While being trained sorely on real human images, our method is empowered with strong domain generalization capability (e.g., Fig. 12) by leveraging the generative prior of a pre-trained stable diffusion model. However, we acknowledge that visual artifacts are possible due to the appearance bias originating from the training data distribution.
@@ Appendix C Limitation and Future Work
While the image coherence is largely strengthened with our cross-view module and 3D-aware generation, we still observe occasionally flickering artifacts in unobserved regions. We leave the exploration of longer-range consistent view manipulation as future work. In this work, the appearance is formulated to be sourced from the reference images only. This could result in some loss of identity given the limited appearance context. In the future, we would like to extend our framework such that the identity can be multi-sourced from e.g., text and personalized Loras [23]. As discussed above, we also include visualizations of failure cases in Figure 15. Rows (a) through (f) display artifacts in areas not observed, accompanied by changes in identity, particularly noticeable in (b), (c), and (d). In cases (a), (e), and (f), it is evident that the model struggles to accurately replicate secondary elements from the reference image, such as sunglasses in (f), hands and flowers in (e), and leaves in (a), leading to inconsistent outcomes. One potential issue we’ve identified stems from the use of a 2D diffusion backbone that integrates 3D-Aware information. This approach, while innovative, may lead to minor inconsistencies and deviations, especially in challenging depictions. Addressing these limitations is an important area that should be addressed in future work.
Appendix D Ethic Consideration
We acknowledge the profound capabilities of the diffusion model as a powerful generative model. The framework proposed in our paper could, theoretically, be utilized to compromise multi-perspective facial recognition systems. We assert that the model and the accompanying research code are intended exclusively for advancing scientific research and must not be used for illicit purposes.